Can you help me with this question please.
Data on college graduates' employment indicate that 40% of the graduates find a part-time job within a year of their graduation, 35% of them find a full-time job and the rest of them are unemployed.
Round your final answer to 1 decimal place and express your answer in percent form (i.e. 30.0% instead of 0.3)
If we randomly select 4 graduate students within a year of their graduation, what is the probability that at least one of them has a job (full-time or part-time)?

Answer by ikleyn(52771)

.
Data on college graduates' employment indicate that 40% of the graduates find a part-time job within a year of their graduation,
35% of them find a full-time job and the rest of them are unemployed.
Round your final answer to 1 decimal place and express your answer in percent form (i.e. 30.0% instead of 0.3)
If we randomly select 4 graduate students within a year of their graduation,
what is the probability that at least one of them has a job (full-time or part-time)?
~~~~~~~~~~~~~~~~~


The probability that one single college graduate student has a job  within a year of the graduation is  0.4 + 0.35 = 0.75.

The probability that one single college graduate student is unemployed  within a year of the graduation is  
the complement to it, i.e.  1 - 0.75 = 0.25.


The probability that of 4 randomly selected college graduate students all 4 are unemployed is, therefore,

    P = 0.25^4 = 0.0039


The probability that of 4 randomly selected college graduate students at least one has has a job is the COMPLEMENT to it, i.e.

    1 - 0.25^4 = 1 - 0.0039 = 0.996 = 99.6%.    ANSWER

Solved, answered and explained.